OpenStudy (anonymous):

scientific notation simplify .00342x10^-5

OpenStudy (ash2326):

@Diana97 did you try to solve this?

OpenStudy (anonymous):

yea

OpenStudy (ash2326):

what did you get?

OpenStudy (anonymous):

i got 3.42000=3.42x10^6

OpenStudy (ash2326):

we have \[\large 0.00342 \times 10^{-5}\] notice that we already have a '-' sign with the 10 power now I'll take the decimal point to one step right doing so will decrease my power by 1 \[\large 0.00342 \times 10^{-5}=00.0342\times 10^{-5-1}\] do you get this?

OpenStudy (anonymous):

yea almost

OpenStudy (anonymous):

exponent will be 4

OpenStudy (ash2326):

it'll be -6 \[\large -5-1=-6\]

OpenStudy (anonymous):

oh ok

OpenStudy (ash2326):

it's not complete yet, shift once more the decimal point to the right what do you get?